OpenStudy (anonymous):

@blurbendy Walt and Amy each have the same number of pennies. Walt made equal stacks of 10 pennies with none left over. Amy made equal stacks of 8 pennies with none left over. What is the least number of pennies that each person could have?

OpenStudy (blurbendy):

80 pennies 10*8 = 80

OpenStudy (anonymous):

Okay, thanks! c:

OpenStudy (anonymous):

Wait, could you explain that?

OpenStudy (anonymous):

Like, with sentences.

OpenStudy (blurbendy):

Sorry, I made a mistake

OpenStudy (blurbendy):

The answer is 40

OpenStudy (blurbendy):

5*8=40; 4*10=40 5 equal stacks of 8 pennies with none left over = 40 4 equal stacks of 10 pennies = 40 So, the least number is 40 pennies.

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (blurbendy):

yup